A firm employs 20 workers at a wage rate of $20 per hour, and 25 units of capital at a cost of $30 per unit per hour. The marginal product of labor is X, and the marginal product of capital is 3. The firm is producing efficiently given what value of X?
Answer #1

At the optimal combination of inputs : w/r = MPL/MPK

Here , w= $20 per hour

r = $30 per unit

MPL = X

MPK = 3

20/30 = X/3

X = 2

This implies that the firm is producing efficiently given the value of X that is MPL = 2.